# Cosmic (super)strings with a time-varying tension

### v1

@@ 2.1.2 Warped throat
The second possibility is to localise the strings at the bottom of warped throat geometry attached to a “bulk” Calabi-Yau geometry, so that the original string tension is redshifted by an amount which can be tuned exponentially, depending on the geometric parameters. In particular, the latter can be described by a metric of the form Dasgupta:1999ss ; Giddings:2001yu
$ds^{2}=e^{2A(y)}g_{\mu\nu}dx^{\mu}dx^{\nu}+e^{-2A(y)}g_{mn}dx^{n}dx^{m},$ (3)
@@
$m_{s}=e^{2A(y_{0})}\frac{g_{s}^{1/4}M_{P}}{\sqrt{4\pi\mathcal{V}}},$ (4)
where the volume is not assumed to be exponentially large (although one must still work in the large volume limit to apply the supergravity approximation). The interest here lies in the fact that brane-antibrane inflation relies on a deep throat to obtain slow-roll (see Baumann:2014nda for extensive discussion on the topic, and Burgess:2022nbx ; Cicoli:2024bwq for recent work). This is also a crucial ingredient in the one of the most well known (and debated) constructions of de Sitter vacua in String Theory, known as KKLT Kachru:2003aw . In particular, a string theoretic embedding of brane inflation within string compactifications in the KKLT framework was achieved in Kachru:2003sx . The vanilla KKLT does not easily allow for modifications to the standard cosmological history and a time-dependent string tension through a rolling volume. This is simply because there is not enough space for the volume to roll before it hits the minimum of its potential. Although we will not pursue this further, let us however notice that there is in principle a different possibility for a time-dependent tension in this case. There is a new modulus, the conifold, which determines the magnitude of the warping factor [^9] Post-inflationary dynamics of this modulus leading to a Hagedorn phase were considered in Frey:2005jk .
@@ 2.2.1 Local $U(1)$ strings
Let us consider the simple example of a $U(1)$ gauge theory which is spontaneously broken. The tension is given by the VEV of the Higgs field that breaks the symmetry in the vacuum. This VEV, however, is an energy scale which in string theory must either be the fundamental string scale or arise dynamically. To see this, focusing in IIB compactifications (where gauge fields arise as open strings stretching between branes - see Ibanez:2012zz ; Marchesano:2022qbx ; Marchesano:2024gul for reviews), notice that the relevant part of the action for the worldvolume theory of a Dp-brane has the schematic form
$S_{\text{DBI}}=T_{p}\int{d^{p+1}x\,\sqrt{h}\left(-\frac{1}{4}F_{AB}F^{AB}-D_{\mu}\phi D^{\mu}\phi^{*}-V(\phi,\phi^{*})\right)}\,.$ (5)
Here $T_{p}\sim\alpha^{\prime-(p+1)/2}$ and $h_{AB}$ is the pullback of the ambient metric (3) on the worldvolume of the brane. To illustrate the point, let us consider a toy model D3-brane with a worldvolume abelian-Higgs model:
$V(\phi,\phi^{*})=\frac{\lambda}{4}\left(|\phi|^{2}-v^{2}\right)^{2}\,,$ (6)
@@
$S_{cn}=\int{d^{3}x\sqrt{h}\left(-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-D_{\mu}\tilde{\phi}D^{\mu}\tilde{\phi}^{*}-\frac{\tilde{\lambda}}{4}\left(\tilde{\phi}^{2}-\tilde{v}^{2}\right)^{2}\right)}\,.$ (8)
@@
$\tilde{\phi}=\phi\sqrt{T_{3}}\,\qquad\tilde{\lambda}=\lambda/T_{3}\,,\qquad\tilde{v}=v\,\sqrt{T_{3}}\,,$ (9)
(the gauge field also needs to be redefined but this is not important for the present discussion). The tension of the string is therefore given by
$\mu=2\pi\tilde{v}^{2}=\frac{(v\,\cdot M_{p})^{2}}{4\mathcal{V}^{2}}M_{p}^{2}\,.$ (10)
We thus conclude that the tension of the string (in Planck units) generically features a dependence on the volume modulus of the compactification, inherited from the dependence of the Planck scale in this modulus. In realistic scenarios, this scalar must acquire a mass and develop a constant vacuum expectation value, at least, before Big Bang Nucleosynthesis, since time-variations of the Planck scale from this time are very constrained (see the discussion in Cicoli:2023opf ). However, in concrete constructions these moduli are generically very light with respect to the string scale and so it is plausible that they might have been cosmologically active.
@@ 2.2.2 Axion strings
$T_{IR}\simeq\pi f_{a}^{2}\log\left(\frac{r_{\text{IR}}}{r_{\text{core}}}\right),$ (11)
where $r_{\text{core}}$ is the core radius (of order $m_{s}^{-1}$ in string theory examples) and $r_{\text{IR}}$ an IR cutoff, usually taken to be the distance to the next string in the network. The second contribution comes from the string core, which is a singular object from the EFT perspective, and carries information about the UV completion (see Lanza:2020qmt ; Lanza:2021udy ). In the case of string theory axions, the core singularity is resolved in terms of fundamental strings or branes wrapping an internal cycle of the compactification manifold. Therefore, the tension will be equal to the product of a $p+1-$brane tension and the volume of the $p-$cycle it is wrapping in the extra dimensions:
$T_{UV}=T_{p+1}\,{\rm{Vol}\Sigma_{p}}=\frac{g_{s}^{\frac{p-3}{4}}}{2^{\frac{p+1}{2}}(2\pi)^{\frac{p-1}{2}}}\frac{M_{P}^{p+1}}{\mathcal{V}^{\frac{p+1}{2}}}{\rm{Vol}\Sigma_{p}}$ (12)
@@
$T=s+ia.$ (13)
In string theory compactifications, axions arise from the dimensional reduction of higher forms, and inherit a discrete shift symmetry from the higher dimensional gauge symmetry Witten:1984dg ; Svrcek:2006yi . As an example, one may consider $C_{4}$ axions Conlon:2006tq ; Cicoli:2012sz in type IIB flux compactifications, obtained by integrating the $C_{4}$ form on 4-cycles $\Sigma_{i}$ of the compactification manifold:
$c_{i}=\frac{1}{(2\pi)^{2}\alpha^{\prime 2}}\int_{\Sigma_{i}}C_{4}.$ (14)
@@
$\mathcal{L}\supset\frac{k^{2}}{4}\frac{\partial_{\mu}\tau\partial^{\mu}\tau}{\tau^{2}}+\frac{k^{2}}{4}\frac{\partial_{\mu}a\partial^{\mu}a}{\tau^{2}},$ (17)
@@
$f_{a}=M_{P}\,e^{-\frac{\sqrt{2}}{k}\langle\varphi\rangle},\quad\quad\text{where}\quad\varphi=\frac{k}{\sqrt{2}}\log\tau.$ (18)
It follows that both contributions to the axion string tension, $T_{IR}$ and $T_{UV}$, can vary with the moduli. Indeed, it was recently argued in Reece:2024wrn (See also Benabou:2023npn ) that the core tension of a fundamental axion string can be estimated from the decay constant as [^11] In any case, the tension must satisfy the bound $T_{UV}\leq 2\pi M_{p}f_{a}$ to be consistent with the axion Weak Gravity Conjecture (aWGC) Arkani-Hamed:2006emk .
$T_{UV}=2\pi S_{\rm{inst}}f_{a}^{2}$ (19)
in a variety of examples (reproducing e.g. the scalings in (12)), where $S_{\rm{inst}}$ is the action of the instanton generating the axion potential. We thus conclude that both superstrings as well as field theory (global and local) strings inherit dependence in moduli in a stringy setting, which potentially translates to (space)time dependence if the latter have non-trivial profiles.
@@ 3 String networks with constant tension
In this section, we introduce the scaling regime and outline the effects of a reduced inter-commutation probability in the energy density of a scaling cosmic string network with constant tension. This section is mostly review material but will be useful later in section 4 where we will generalize the scenario. It is well known that, modulo the subtleties discussed in the introduction, cosmic string networks reach a configuration dubbed as the scaling regime Kibble:1984hp ; Bennett:1985qt ; Austin:1993rg (see also Vilenkin:2000jqa for a pedagogical discussion): a self-similar configuration in which the energy density $\rho$ of the system remains a fixed fraction of the energy density of the background (which we implicitly assume is much larger than that of the strings, thus leading the expansion of the Universe).
@@ 3.1 Equations of motion
$\gamma_{ab}=\frac{\partial X^{\mu}(\zeta)}{\partial\zeta^{a}}\frac{\partial X^{\nu}(\zeta)}{\partial\zeta^{b}}g_{\mu\nu},\quad\quad\text{where}\quad\quad g_{\mu\nu}=a^{2}(\tau)\eta_{\mu\nu}$ (20)
@@
$S=-\mu\int d^{2}\zeta\,\sqrt{-\gamma}.$ (21)
@@
$\frac{1}{\sqrt{-\gamma}}\partial_{a}\left(\sqrt{-\gamma}\gamma^{ab}g_{\rho\sigma}X^{\sigma}_{\,,b}\right)-\frac{1}{2}\partial_{\rho}g_{\lambda\sigma}\gamma^{ab}X^{\lambda}_{\,,a}X^{\sigma}_{\,,b}=0,$ (22)
@@
$\varepsilon\equiv\sqrt{\frac{x^{\prime 2}}{1-\dot{x}^{2}}},$ (23)
the equation of motion (22) can be recast as
$\dot{\epsilon}+2\frac{\dot{a}}{a}\varepsilon\dot{x}^{2}=0$ (24)
@@
$\ddot{x}_{i}+2\frac{\dot{a}}{a}(1-\dot{x}^{2})\dot{x}_{i}-\frac{1}{\varepsilon}\left(\frac{x_{i}^{{}^{\prime}}}{\varepsilon}\right)^{{}^{\prime}}=0.$ (25)
@@ 3.2 VOS model
The scaling behaviour mentioned in the first paragraph is well captured by the VOS model  Martins:1996jp ; Martins:2000cs , which describes the evolution of a string network in terms of a characteristic length $L$, defined through
$\rho=\frac{\mu}{L^{2}}\,,$ (27)
where $\mu$ is the string tension and $v$ the rms string velocity defined above. $L$ also admits an interpretation as the average distance between strings, and is sometimes called the correlation length. The two fundamental equations in the VOS model are essentially integrated versions of the microscopic equations of motion for an ideal (Nambu-Goto) string, (24)- (25), supplemented with a loop-chopping parameter $\tilde{c}$ which captures the energy loss of the network into loops. The equations read
$\frac{{\rm d}v}{{\rm d}t}=(1-v^{2})\left[\frac{k(v)}{L}-2Hv\right]$ (28)
@@
$L=\xi t\,\qquad\quad\xi=\sqrt{\frac{n}{8}\frac{k(\bar{v})(k(\bar{v})+\tilde{c})}{1-\frac{2}{n}}},$ (30)
where $n$ characterises the cosmological period through $\rho\sim a^{-n}$ (i.e: $a(t)\sim t^{2/n}$) and the important quantity $c$ quantifies the efficiency of energy loss of the network (more on this later). Following Cui:2018rwi , the root mean squared velocity $\bar{v}$ is given, for our purposes, by the expression
$\bar{v}=\sqrt{\frac{n}{2}\frac{k(\bar{v})}{k(\bar{v})+\tilde{c}}\left(1-\frac{2}{n}\right)}\,.$ (31)
@@
$\left\langle\left(1-\dot{\mathbf{x}}^{2}\right)(\dot{\mathbf{x}}\cdot\hat{\mathbf{u}})\right\rangle\equiv kv\left(1-v^{2}\right),$ (32)
where $\hat{\mathbf{u}}$ is the unit curvature vector of the string. It is known in the literature as the momentum parameter, and encodes information about small scale structure of the string. A phenomenologically motivated expression is Martins:2000cs
$k(v)=\frac{2\sqrt{2}}{\pi}\left(1-v^{2}\right)\left(1+2\sqrt{2}v^{3}\right)\frac{1-8v^{6}}{1+8v^{6}},$ (33)
which reproduces the correct non-relativistic limit and vanishes for the flat spacetime value of $v^{2}=\frac{1}{2}$. For an expanding spacetime, the rms string velocity will asymptote to a value just below $v^{2}<\frac{1}{2}$ilenkin:2000jqa , so that $k(\bar{v})$ is small and positive in accordance with (30)-(31). In upcoming sections we will argue that if the tension decreases sufficiently fast with respect to the expansion of the Universe, a scaling regime is achieved with $\bar{v}^{2}>1/2$ and $k(\bar{v})<0$. Before concluding the section, we note that the scaling behaviour has been also studied in the context of cosmic superstrings Sakellariadou:2004wq ; Avgoustidis:2005nv ; Pourtsidou:2010gu ; Avelino:2012qy ; Sousa:2013aaa ; Sousa:2016ggw , where the reconnection probability $P$ is smaller than one. The qualitative features of the discussion are not altered and a scaling regime is achieved, albeit the different quantities ($c,\xi$, etc) inherit a $P$-dependence. In the following we will simply assume the VOS model.

### v2

@@ 2.1.2 Warped throat
The second possibility is to localise the strings at the bottom of warped throat geometry attached to a “bulk” Calabi-Yau geometry, so that the original string tension is redshifted by an amount which can be tuned exponentially, depending on the geometric parameters. In particular, the latter can be described by a metric of the form [73, 63]
$ds^{2}=e^{2A(y)}g_{\mu\nu}dx^{\mu}dx^{\nu}+e^{-2A(y)}g_{mn}dy^{n}dy^{m},$ (2.3)
@@
$\mu=m_{s}^{2}=e^{4A(y_{0})}\frac{g_{s}^{1/2}M_{P}^{2}}{4\pi\mathcal{V}},$ (2.4)
where the volume is not assumed to be exponentially large (although one must still work in the large volume limit to apply the supergravity approximation). The interest here lies in the fact that brane-antibrane inflation relies on a deep throat to obtain slow-roll (see [74] for extensive discussion on the topic, and [51, 52] for recent work). This is also a crucial ingredient in the one of the most well known (and debated) constructions of de Sitter vacua in String Theory, known as KKLT [75]. In particular, a string theoretic embedding of brane inflation within string compactifications in the KKLT framework was achieved in [50]. The vanilla KKLT does not easily allow for modifications to the standard cosmological history and a time-dependent string tension through a rolling volume. This is simply because there is not enough space for the volume to roll before it hits the minimum of its potential. Although we will not pursue this further, let us however notice that there is in principle a different possibility for a time-dependent tension in this case. There is a new modulus, the conifold, which determines the magnitude of the warping factor. [^11] Post-inflationary dynamics of this modulus leading to a Hagedorn phase were considered in [76].
@@ 2.2.1 Local $U(1)$ strings
Let us consider the simple example of a $U(1)$ gauge theory which is spontaneously broken. The tension is given by the VEV of the Higgs field that breaks the symmetry in the vacuum. This VEV, however, is an energy scale which in string theory must either be the fundamental string scale or arise dynamically. To see this, focusing in IIB compactifications (where gauge fields arise as open strings stretching between branes - see [77, 78, 79] for reviews), notice that the relevant part of the action for the worldvolume theory of a Dp-brane has the schematic form
$S_{\text{DBI}}=T_{p}\int{d^{p+1}x\,\sqrt{h}\left(-\frac{1}{4}F_{AB}F^{AB}-D_{\mu}\phi D^{\mu}\phi^{*}-V(\phi,\phi^{*})\right)}\,.$ (2.5)
Here $T_{p}\sim\alpha^{\prime-(p+1)/2}$ and $h_{AB}$ is the pullback of the ambient metric (2.3) on the worldvolume of the brane. To illustrate the point, let us consider a toy model D3-brane with a worldvolume abelian-Higgs model:
$V(\phi,\phi^{*})=\frac{\lambda}{4}\left(|\phi|^{2}-v^{2}\right)^{2}\,,$ (2.6)
@@
$S_{cn}=\int{d^{3}x\sqrt{h}\left(-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-D_{\mu}\tilde{\phi}D^{\mu}\tilde{\phi}^{*}-\frac{\tilde{\lambda}}{4}\left(\tilde{\phi}^{2}-\tilde{v}^{2}\right)^{2}\right)}\,.$ (2.8)
@@
$\tilde{\phi}=\phi\sqrt{T_{3}}\,\qquad\tilde{\lambda}=\lambda/T_{3}\,,\qquad\tilde{v}=v\,\sqrt{T_{3}}\,,$ (2.9)
(the gauge field also needs to be redefined but this is not important for the present discussion). The effect of this field redefinition is to introduce a moduli dependence on the scalar field vev, which is further inherited by the string tension. In particular, the latter is given by
$\mu=2\pi\tilde{v}^{2}=\frac{(v\,\cdot M_{p})^{2}}{4\mathcal{V}^{2}}M_{p}^{2}\,.$ (2.10)
We thus conclude that the tension of the string (in Planck units) generically features a dependence on the volume modulus of the compactification, inherited from the dependence of the Planck scale in this modulus. In realistic scenarios, this scalar must acquire a mass and develop a constant vacuum expectation value, at least, before Big Bang Nucleosynthesis, since time-variations of the Planck scale from this time are very constrained (see the discussion in [28]). However, in concrete constructions these moduli are generically very light with respect to the string scale and so it is plausible that they might have been cosmologically active.
@@ 2.2.2 Axion strings
$T_{IR}\simeq\pi f_{a}^{2}\log\left(\frac{r_{\text{IR}}}{r_{\text{core}}}\right),$ (2.11)
where $r_{\text{core}}$ is the core radius (of order $m_{s}^{-1}$ in string theory examples) and $r_{\text{IR}}$ an IR cutoff, usually taken to be the distance to the next string in the network. The second contribution comes from the string core, which is a singular object from the EFT perspective, and carries information about the UV completion (see [80, 81]). In the case of string theory axions, the core singularity is resolved in terms of fundamental strings or branes wrapping an internal cycle of the compactification manifold. Therefore, the tension will be equal to the product of a $p+1-$brane tension and the volume of the $p-$cycle it is wrapping in the extra dimensions:
$T_{UV}=T_{p+1}\,{\rm{Vol}\Sigma_{p}}=\frac{g_{s}^{\frac{p-3}{4}}}{2^{\frac{p+1}{2}}(2\pi)^{\frac{p-1}{2}}}\frac{M_{P}^{p+1}}{\mathcal{V}^{\frac{p+1}{2}}}{\rm{Vol}\Sigma_{p}}$ (2.12)
@@
$T=s+ia.$ (2.13)
In string theory compactifications, axions arise from the dimensional reduction of higher forms, and inherit a discrete shift symmetry from the higher dimensional gauge symmetry [82, 83]. As an example, one may consider $C_{4}$ axions [84, 85] in type IIB flux compactifications, obtained by integrating the $C_{4}$ form on 4-cycles $\Sigma_{i}$ of the compactification manifold:
$c_{i}=\frac{1}{(2\pi)^{2}\alpha^{\prime 2}}\int_{\Sigma_{i}}C_{4}.$ (2.14)
@@
$\mathcal{L}\supset\frac{k^{2}}{4}\frac{\partial_{\mu}\tau\partial^{\mu}\tau}{\tau^{2}}+\frac{k^{2}}{4}\frac{\partial_{\mu}a\partial^{\mu}a}{\tau^{2}},$ (2.17)
@@
$f_{a}=M_{P}\,e^{-\frac{\sqrt{2}}{k}\langle\varphi\rangle},\quad\quad\text{where}\quad\varphi=\frac{k}{\sqrt{2}}\log\tau.$ (2.18)
It follows that both contributions to the axion string tension, $T_{IR}$ and $T_{UV}$, can vary with the moduli. Indeed, it was recently argued in [86] (See also [87]) that the core tension of a fundamental axion string can be estimated from the decay constant as [^13] In any case, the tension must satisfy the bound $T_{UV}\leq 2\pi M_{p}f_{a}$ to be consistent with the axion Weak Gravity Conjecture (aWGC) [88].
$T_{UV}=2\pi S_{\rm{inst}}f_{a}^{2}$ (2.19)
in a variety of examples (reproducing e.g. the scalings in (2.12)), where $S_{\rm{inst}}$ is the action of the instanton generating the axion potential. We thus conclude that both superstrings as well as field theory (global and local) strings inherit dependence in moduli in a stringy setting, which potentially translates to (space)time dependence if the latter have non-trivial profiles.
@@ 3 String networks with constant tension
In this section, we introduce the scaling regime and outline the effects of a reduced inter-commutation probability in the energy density of a scaling cosmic string network with constant tension. This section is mostly review material but will be useful later in section 4 where we will generalize the scenario. It is well known that, modulo the subtleties discussed in the introduction, cosmic string networks reach a configuration dubbed as the scaling regime [2, 89, 90] (see also [4] for a pedagogical discussion): a self-similar configuration in which the energy density $\rho$ of the system remains a fixed fraction of the energy density of the background (which we implicitly assume is much larger than that of the strings, thus leading the expansion of the Universe).
@@ 3.1 Equations of motion
$\gamma_{ab}=\frac{\partial X^{\mu}(\zeta)}{\partial\zeta^{a}}\frac{\partial X^{\nu}(\zeta)}{\partial\zeta^{b}}g_{\mu\nu},\quad\quad\text{where}\quad\quad g_{\mu\nu}=a^{2}(\tau)\eta_{\mu\nu}$ (3.1)
@@
$S=-\mu\int d^{2}\zeta\,\sqrt{-\gamma}.$ (3.2)
@@
$\frac{1}{\sqrt{-\gamma}}\partial_{a}\left(\sqrt{-\gamma}\gamma^{ab}g_{\rho\sigma}X^{\sigma}_{\,,b}\right)-\frac{1}{2}\partial_{\rho}g_{\lambda\sigma}\gamma^{ab}X^{\lambda}_{\,,a}X^{\sigma}_{\,,b}=0,$ (3.3)
@@
$\varepsilon\equiv\sqrt{\frac{x^{\prime 2}}{1-\dot{x}^{2}}},$ (3.4)
the equation of motion (3.3) can be recast as
$\dot{\varepsilon}+2\frac{\dot{a}}{a}\varepsilon\dot{x}^{2}=0$ (3.5)
@@
$\ddot{x}_{i}+2\frac{\dot{a}}{a}(1-\dot{x}^{2})\dot{x}_{i}-\frac{1}{\varepsilon}\left(\frac{x_{i}^{{}^{\prime}}}{\varepsilon}\right)^{{}^{\prime}}=0.$ (3.6)
@@ 3.2 VOS model
The scaling behaviour mentioned in the first paragraph is well captured by the VOS model  [36, 37], which describes the evolution of a string network in terms of a characteristic length $L$, defined through
$\rho=\frac{\mu}{L^{2}}\,,$ (3.8)
where $\mu$ is the string tension and $v$ the rms string velocity defined above. $L$ also admits an interpretation as the average distance between strings, and is sometimes called the correlation length. The two fundamental equations in the VOS model are essentially integrated versions of the microscopic equations of motion for an ideal (Nambu-Goto) string, (3.5)- (3.6), supplemented with a loop-chopping parameter $\tilde{c}$ which captures the energy loss of the network into loops. The equations read
$\frac{{\rm d}v}{{\rm d}t}=(1-v^{2})\left[\frac{k(v)}{L}-2Hv\right]$ (3.9)
@@
$L=\xi t\,\qquad\quad\xi=\sqrt{\frac{n}{8}\frac{k(\bar{v})(k(\bar{v})+\tilde{c})}{1-\frac{2}{n}}},$ (3.11)
where $n$ characterises the cosmological period through $\rho\sim a^{-n}$ (i.e: $a(t)\sim t^{2/n}$) and the important quantity $c$ quantifies the efficiency of energy loss of the network (more on this later). Following [26], the root mean squared velocity $\bar{v}$ is given, for our purposes, by the expression
$\bar{v}=\sqrt{\frac{n}{2}\frac{k(\bar{v})}{k(\bar{v})+\tilde{c}}\left(1-\frac{2}{n}\right)}\,.$ (3.12)
@@
$\left\langle\left(1-\dot{\mathbf{x}}^{2}\right)(\dot{\mathbf{x}}\cdot\hat{\mathbf{u}})\right\rangle\equiv kv\left(1-v^{2}\right),$ (3.13)
where $\hat{\mathbf{u}}$ is the unit curvature vector of the string. It is known in the literature as the momentum parameter, and encodes information about small scale structure of the string. A phenomenologically motivated expression is [37]
$k(v)=\frac{2\sqrt{2}}{\pi}\left(1-v^{2}\right)\left(1+2\sqrt{2}v^{3}\right)\frac{1-8v^{6}}{1+8v^{6}},$ (3.14)
which reproduces the correct non-relativistic limit and vanishes for the flat spacetime value of $v^{2}=\frac{1}{2}$. For an expanding spacetime, the rms string velocity will asymptote to a value just below $v^{2}<\frac{1}{2}$[4], so that $k(\bar{v})$ is small and positive in accordance with (3.11)-(3.12). In upcoming sections we will argue that if the tension decreases sufficiently fast with respect to the expansion of the Universe, a scaling regime is achieved with $\bar{v}^{2}>1/2$ and $k(\bar{v})<0$. Before concluding the section, we note that the scaling behaviour has been also studied in the context of cosmic superstrings [91, 92, 93, 94, 16, 95], where the reconnection probability $P$ is smaller than one. The qualitative features of the discussion are not altered and a scaling regime is achieved, albeit the different quantities ($c,\xi$, etc) inherit a $P$-dependence. In the following we will simply assume the VOS model.